An office supply store sells seven tablets and 4 pens for $6.40. Also, two tablets and 19 pens cost $5.40. Find the price of each.
Answer by Paul(988)

Let the cost of the tablets be x
Let the cost of the pens be y
7x+4y=6.4 --->y=1.6-1.75x (subsitution)
Second equatioN: 2x+19y=5.4
Subsitute for y
2x+19(1.6-1.75x)=5.4
2x-33.25x=5.4-30.4
31.25x=25
x=0.80

y=1.6-1.75(0.8)
y=0.20
Hence, the cost of a single tablet is $0.80 and a cost of a single pen is $0.20.
Paul.